OpenStudy (anonymous):

find the integral of dz/sqr(z^2+r^2)

OpenStudy (anonymous):

pls guys

OpenStudy (jhannybean):

Is Z in respect to r?

OpenStudy (anonymous):

i dont think so

OpenStudy (jhannybean):

\[\int\limits \frac{ dx }{ \sqrt{x^2 \pm a^2}}=\ln|x+\sqrt{x^2 \pm a^2}| \] be of any help? only considering r is a constant.. i might be wrong :(

OpenStudy (dan815):